### v1

@@ Abstract
Nowadays, pre-trained encoders are widely used in medical image segmentation because of their ability to capture complex feature representations. However, the existing models fail to effectively utilize the rich features obtained by the pre-trained encoder, resulting in suboptimal segmentation results. In this work, a novel U-shaped model, called FIF-UNet, is proposed to address the above issue, including three plug-and-play modules. A channel spatial interaction module (CSI) is proposed to obtain informative features by establishing the interaction between encoder stages and corresponding decoder stages. A cascaded conv-SE module (CoSE) is designed to enhance the representation of critical features by adaptively assigning importance weights on different feature channels. A multi-level fusion module (MLF) is proposed to fuse the multi-scale features from the decoder stages, ensuring accurate and robust final segmentation. Comprehensive experiments on the Synapse and ACDC datasets demonstrate that the proposed FIF-UNet outperforms existing state-of-the-art methods, which achieves the highest average DICE of 86.05% and 92.58%, respectively.
@@ 1 Introduction
Medical image segmentation refers to extracting regions of interest from medical images, such as organs, diseased areas, etc., which can help physicians make diagnoses and formulate treatment plans. In medical image analysis, medical image segmentation is a fundamental task that provides accurate anatomical structure information for subsequent image analysis and quantification. Automatic segmentation provides high-precision segmentation masks (size, shape, and location of lesions) to physicians for accurately identifying and segmenting specific structures or lesions in images. Compared to handcrafted segmentation, automatic segmentation can significantly reduce the time and cost of processing medical images and improve the efficiency of data processing. In addition, automatic segmentation tools can mitigate the influence of human-related subjective factors, providing higher accuracy and consistency, which is particularly important for large-scale medical image datasets.
@@
Although the UNet models demonstrate the desired performance in medical image segmentation tasks, they still cannot capture global contextual information due to the limited receptive field. To address this issue, the Transformer blocks (Vaswani et al. (2017)) are incorporated into UNet architectures to enhance the global feature integration and contextual understanding, such as SwinUNet (Cao et al. (2022)), TransUNet (Chen et al. (2021)), MISSFormer (Huang et al. (2021)), UNRTR (Hatamizadeh et al. (2022)), and so on. However, the Transformer lacks inductive biases in CNNs, such as translation invariance and local feature learning ability, which makes it hard to achieve the expected performance with insufficient training samples. In the pre-training Vision Transformer (ViT) (Dosovitskiy et al. (2020)), the backbone network is pre-trained on large-scale common datasets and further fine-tuned with limited samples for certain tasks, which demonstrates significant efficacy in practical applications and reduces the requirements for labeled samples, such as Efficientnet (Tan and Le (2019)), ConvNeXt (Liu et al. (2022)), DeepViT (Zhou et al. (2021)), etc. Thanks to recent advances in pre-training strategy, the pre-trained encoders are empowered with rich and generalized feature representations. The key to improving task performance is to design efficient feature interaction and fusion mechanisms in the skip connections and decoder, which has the ability to leverage the semantic information learned by the pre-trained encoder to predict segmentation masks.
To this end, a novel U-shaped model, called FIF-UNet, is proposed to effectively utilize different levels of semantic information by feature interaction and fusion. Considering the semantic gap between the features learned by the different encoder and decoder stages, the feature fusion based on concatenation or element-wise addition operations in the vanilla skip connection inevitably results in inaccurate features and information loss, impacting the learning and generalization ability of the model to support the segmentation task. In this work, a channel spatial interaction (CSI) module, including the cascaded channel interaction unit (CIU) and spatial interaction unit (SIU), is proposed to enhance the skip connections. The CSI is designed to interactively recalibrate the feature maps by capturing the correlation between different levels of semantic information in a learnable manner.
In the decoder, the squeeze-and-excitation network (SENet) (Hu et al. (2018)) is incorporated into the original CNN blocks, i.e., cascaded conv-SE module (CoSE). The CoSE module leverages channel attention to adaptively reweight the features of different channels by modeling the interdependency among channels. The primary purpose is to efficiently select and integrate the crucial features to highlight the target regions and suppress the irrelevant background regions. In addition, to generate the pixel-wise segmentation masks, the up-sampling operations are applied to the consecutive decoder blocks to recover feature resolution. In this procedure, limited fusion in multi-level features leads to the loss of detailed features. In this work, a multi-level fusion module (MLF) is designed to effectively fuse the learned multi-scale features in different decoder stages by feature interactions among intra- and inter-classes.
Extensive experiments are conducted on the open-source Synapse and ACDC datasets to validate the proposed model. The experimental results demonstrate that the proposed approach outperforms other selective baselines, achieving an average DICE of 86.05% and 92.58%, respectively. Most importantly, the proposed three plug-and-play modules contribute to expected performance improvements and the visualization results indicate a confident location.
@@
(1) A new U-shaped model is proposed to fully utilize the multi-level semantic information of the encoder and decoder by feature interaction and fusion. The results demonstrate the robustness of the FIF-UNet and its potential in practical applications.
(2) The CSI module is proposed to iteratively recalibrate the feature maps by capturing the correlation between the features learned by the different encoder and decoder stages, in which the CIU and SIU are cascaded to implement the feature interactions progressively.
(3) In the decoder, the CoSE module is designed to learn critical features based on the local structure and global context to highlight the target regions by incorporating channel attention into the convolution operations.
(4) The MLF module is proposed to efficiently fuse semantic information of different scales from decoder stages to alleviate the problem of detail loss, helping to obtain an accurate and robust final segmentation.
@@ 2 Related Work
As the core blocks of the UNet, CNN-based models were the dominant methods for various computer vision tasks. The Transformer-based models were also regarded as the mainstream due to their recent advancements across many artificial intelligence tasks. Integrating these modules with the UNet emerges as an enhanced strategy to improve the performance of medical image segmentation. In this section, related works are organized as follows:
@@ 2.1 CNN models
Before the ViT model in 2020, CNN-based UNet models were the dominant approaches in the field of medical image segmentation, which efficiently capture local features through convolution operations. However, the original UNet suffered from limited feature extraction capability and the semantic gap between the encoder and decoder. To address these issues, the encoder or decoder modules of the UNet were improved to enhance feature learning. In DUNet (Jin et al. (2019)), the convolution layers of the original UNet were replaced by the deformable convolution layer to capture intricate features. The Attention Gate (AG) was introduced to automatically focus on target structures with different shapes and sizes by employing a large receptive field and semantic contextual information in Attention U-Net (Oktay et al. (2018)). The inception layers of the Google-Net (Punn and Agarwal (2020)) were applied to automate the selection of the variety of layers in the deep network. However, the mentioned improvements were mainly based on local convolution operations, with only a weak ability to capture global contexts.
Other works focused on adjusting the skip connections to alleviate the semantic gap between the encoder and decoder. The Group Aggregation Bridge module (GAB) in EGE-UNet (Ruan et al. (2023)) effectively fused multi-scale information by grouping low-level features, high-level features, and a mask generated by the decoder at each stage. The densely connected skip connections were designed to aggregate features of different semantic scales in UNet++ (Zhou et al. (2019)), resulting in forming a highly flexible feature fusion scheme. A two-round fusion module (i.e., top to bottom and bottom to top) in the skip connections was performed to reduce the semantic gap in FusionU-Net (Li et al. (2024)). However, existing works only utilize the features of different encoder layers to alleviate the semantic gap, neglecting the importance of effective semantic fusion between the encoder layer and the decoder layer.
@@ 2.2 Vision Transformer models
Transformer was initially proposed for natural language processing and opened up new avenues for innovation in computer vision tasks (Dosovitskiy et al. (2020)). The Transformer block allows each element in the input sequence to focus on all other elements by self-attention mechanism, thus constructing pure Transformer models to effectively adapt to complex image scenes and objects of various sizes over CNNs. For example, the hierarchical Swin Transformer with shifted windows was used as a base block to learn global and distant semantic interactions in SwinUNet (Cao et al. (2022)). Gating mechanisms were added to the axial-attention to learn relative positional coding, to further accurately encode long-range interactions in MedT (Valanarasu et al. (2021)). In MISSFormer (Huang et al. (2021)), efficient self-attention and enhanced mix-FFN were introduced to construct an enhanced Transformer block for aligning features with higher consistency. Inspired by dilated convolutions, a dilated Transformer was proposed to perform global self-attention in a dilated manner in D-Former (Wu et al. (2023)), which expanded the receptive field and reduced computational cost without adding patch blocks. However, compared with CNNs, the pure Transformer model is limited by the learning of local features, which impacts the accurate capture of detailed features, especially in the delicate medical image segmentation task.
@@ 2.3 Hybrid CNN-Transformer models
Figure 1: (a) The architecture of FIF-UNet, (b) CoSE Module, (c) UpConv Module, (d) Squeeze-and-Excitation Network, (e) ConvBlock.
@@ 3.1 Overall architecture
The architecture of the FIF-UNet is shown in Figure 1(a), implemented by a U-shaped architecture with symmetric encoder-decoder modules. In the encoder, MaxVit-S (Tu et al. (2022)) serves as the backbone network, which is pre-trained on the ImageNet dataset utilizing an image classification task. Compared to full self-attention, the MaxViT is implemented based on blocked local and dilated global attention to capture both the local and global features, which can be calculated by only linear complexity ($O(n)$, $n$ is the spatial size of an input image). The encoder network consists of 5 stages, including a stem stage and four cascaded MaxViT stages. In the stem stage, two convolution layers are with 96 channels and a kernel size of 3. The stride of the first CNN layer is set to 2 to downsample the input image resolution. The configurations of the MaxViT are with the $\left\{2,2,5,2\right\}$ blocks and generate the feature maps with $\left\{96,192,384,768\right\}$ channels, respectively.
In the skip connections, a CSI module is proposed to dynamically recalibrate the feature maps by the designed CIU and SIU, with the objective of obtaining the informative target features. In the decoder network, each decoder stage is constructed based on the CoSE module and UpConv module. The proposed CoSE module aims to enhance the representation of critical features by incorporating the SENet mechanism into CNNs. The UpConv module upsamples the resolution of the CoSE outputs by bilinear interpolation, followed by a convolution layer to refine the up-sampled feature maps, as in Figure 1(c). Instead of predicting the segmentation tasks based only on the last decoder stage, in this work, an MLF module is innovatively proposed to effectively fuse the outputs of the decoder stages to enhance the segmentation details by integrating intra- and inter-class features.
@@ 3.2 CSI Module
In general, the CSI module is based on the sequential channel interaction unit (CIU) and spatial interaction unit (SIU), as in Figure 1(a). The CSI module is performed on the skip connections of the UNet architecture, by taking feature maps of encoder stage $i$ ($2,3,4$) and decoder stage $i+1$ as inputs. The purpose of the design is listed below:
  • The CIU focuses on achieving semantic alignment between the encoder and decoder features along the channel dimension by interactively adjusting the channel weights.
  • The SIU is designed to capture spatial correlations among pixels, which is expected to enable semantic complementation to support feature learning.
Finally, the output of the CSI module is fed into the decoder stage $i$. It is believed that the CSI can be integrated into any UNet architecture, enabling skip connections to effectively fuse different levels of semantics to obtain informative feature maps.
@@ 3.2.1 CIU
As mentioned before, the inputs of the CIU are the feature maps of encoder stage $i$ ($X$) and decoder stage $i+1$ ($Y$). The core idea of the CIU is to adjust the importance of input features ($X$ or $Y$) based on another counterpart feature map ($Y$ or $X$), implemented by parallel paths for the inputs separately (Figure 2(a)).
@@ 3.2.2 SIU
In the right part of the X-shaped path, the sigmoid activation function is performed to project the weight elements to $[0,1]$, which is further performed on the SIU inputs to reweight their pixel-wise importance. Finally, the output feature map is formulated by addition operations, as in:
@@
In summary, the interaction and fusion operations consider the correlation among pixels in input feature maps and reassign their importance weights to achieve semantic alignment and complementation. By cascading the CIU and SIU modules, the CSI module is expected to enhance the skip connections of the UNet architecture to deliver rich features to the decoder.

### v2

@@ Abstract
Nowadays, pre-trained encoders are widely used in medical image segmentation due to their strong capability in extracting rich and generalized feature representations. However, existing methods often fail to fully leverage these features, limiting segmentation performance. In this work, a novel U-shaped model is proposed to address the above issue, including three plug-and-play modules. A channel spatial interaction module is introduced to improve the quality of skip connection features by modeling inter-stage interactions between the encoder and decoder. A channel attention-based module integrating squeeze-and-excitation mechanisms with convolutional layers is employed in the decoder blocks to strengthen the representation of critical features while suppressing irrelevant ones. A multi-level fusion module is designed to aggregate multi-scale decoder features, improving spatial detail and consistency in the final prediction. Comprehensive experiments on the synapse multi-organ segmentation dataset and automated cardiac diagnosis challenge dataset demonstrate that the proposed model outperforms existing state-of-the-art methods, achieving the highest average Dice score of 86.05% and 92.58%, yielding improvements of 1.15% and 0.26%, respectively. In addition, the proposed model provides a balance between accuracy and computational complexity, with only 86.91 million parameters and 23.26 giga floating-point operations.
@@ 1 Introduction
Organ segmentation in medical images plays a vital role in the field of healthcare, providing essential structural information for diagnosis, treatment planning, and clinical decision-making. Accurate delineation of organs is fundamental for downstream tasks, such as radiation therapy planning, surgical navigation, and quantitative assessment of organ morphology and function. Traditionally, organ segmentation has been performed manually by expert clinicians, which is time-consuming and subject to variability due to human error and inter-observer differences. In contrast, automatic organ segmentation methods can substantially reduce the time and effort required for image analysis, thereby improving clinical efficiency. Moreover, by minimizing human-related subjectivity, these methods offer enhanced consistency, reproducibility, and accuracy, which are especially critical in high-stakes clinical environments.
@@
Although the UNet models demonstrate the desired performance in medical image segmentation tasks, they still cannot capture global contextual information due to the limited receptive field. To address this issue, the Transformer blocks (Vaswani et al. (2017)) are incorporated into UNet architectures to enhance the global feature integration and contextual understanding, such as SwinUNet (Cao et al. (2022)), TransUNet (Chen et al. (2021)), MISSFormer (Huang et al. (2021)), UNRTR (Hatamizadeh et al. (2022)), and so on. However, the Transformer lacks inductive biases in CNNs, such as translation invariance and local feature learning ability, which makes it hard to achieve the expected performance with insufficient training samples. Recent advances in the pre-training Vision Transformer (ViT) (Dosovitskiy et al. (2020)) (e.g., Efficientnet (Tan and Le (2019)), ConvNeXt (Liu et al. (2022)), DeepViT (Zhou et al. (2021)) have enabled the extraction of rich and generalizable visual features from large-scale datasets. These pre-trained encoders significantly reduce the reliance on task-specific labeled data and offer strong performance for downstream tasks. However, efficiently leveraging such powerful encoder features in the decoder remains a critical challenge, particularly for pixel-wise prediction tasks like medical image segmentation.
To address this, we propose a novel U-shaped model (FIF-UNet) to fully utilize multi-level semantic features extracted by the encoder. The proposed model integrates three purpose-designed components to enhance the information flow between encoder and decoder in a complementary manner: (1) Channel Spatial Interaction (CSI) module, (2) Cascaded convolution-SE (CoSE) module, and (3) Multi-Level Fusion (MLF) module.
The CSI module is designed to explicitly address the semantic gap between encoder and decoder features, cascading a channel interaction unit (CIU) and a spatial interaction unit (SIU) to refine skip connection features. This design enables the model to learn correlations between semantic representations at different levels, thereby delivering more accurate and informative feature maps to the decoder. In the decoder path, the CoSE module is proposed to enhance feature selection by integrating channel attention through Squeeze-and-Excitation (SE) blocks (Hu et al. (2018)) into convolution operations. The CoSE enables the model to adaptively emphasize task-relevant features while suppressing background noise, which is crucial for delineating complex anatomical structures. In addition, to mitigate the detail degradation caused by upsampling operations, the MLF module is proposed to aggregate semantic features across decoder layers. Unlike standard progressive upsampling, MLF performs feature interactions across multiple scales to retain fine-grained information and contextual coherence in the predicted segmentation maps.
In the proposed model, these three components form a coherent feature interaction and fusion strategy to enhance both the representational capacity and generalization ability. Extensive experiments on Synapse and ACDC datasets show that the proposed model outperforms selective competitive baselines, achieving average Dice scores of 86.05% and 92.58%, respectively. The ablation results further confirm the dedicated performance contributions of CSI, CoSE, and MLF.
@@
(1) A new U-shaped model, FIF-UNet, is proposed to improve feature exploration from pre-trained encoders through explicit feature interaction and fusion strategies in the decoding process.
(2) The CSI module is designed to enhance the quality of features passed through skip connection, thus providing the decoder with more accurate and semantically enriched feature maps.
(3) The CoSE module is integrated into the decoder to enhance discriminative feature learning by combining convolutional operations with channel attention mechanisms.
(4) The MLF module is introduced to efficiently aggregate multi-scale decoder features, improving the model’s ability to recover detailed structures in segmentation outputs.
@@ 2 Related Work
As the core blocks of the UNet, CNN-based models were the dominant methods for various computer vision tasks. The Transformer-based models were also regarded as mainstream due to their recent advancements across many artificial intelligence tasks. Integrating these modules with the UNet emerges as an enhanced strategy to improve the performance of medical image segmentation. In this section, related works are organized as follows:
@@ 2.1 CNN models
Before the ViT model in 2020, CNN-based UNet models were the dominant approaches in the field of medical image segmentation, which efficiently capture local features through convolution operations. However, the original UNet suffered from limited feature extraction capability and the semantic gap between the encoder and decoder. To address these issues, the encoder or decoder modules of the UNet were improved to enhance feature learning. In DUNet (Jin et al. (2019)), the convolution layers of the original UNet were replaced by the deformable convolution layer to capture intricate features. The inception layers of the Google-Net (Punn and Agarwal (2020)) were applied to automate the selection of the variety of layers in the deep network. However, the mentioned improvements were mainly based on local convolution operations, with only a weak ability to capture global contexts.
Other works focused on adjusting the skip connections to alleviate the semantic gap between the encoder and decoder. The Group Aggregation Bridge module (GAB) in EGE-UNet (Ruan et al. (2023)) effectively fused multi-scale information by grouping low-level features, high-level features, and a mask generated by the decoder at each stage. The densely connected skip connections were designed to aggregate features of different semantic scales in UNet++ (Zhou et al. (2019)), resulting in forming a highly flexible feature fusion scheme. The Attention Gate (AG) was introduced to automatically focus on target structures with different shapes and sizes by employing a large receptive field and semantic contextual information in Attention U-Net (Oktay et al. (2018)). A two-round fusion module (i.e., top to bottom and bottom to top) in the skip connections was performed to reduce the semantic gap in FusionU-Net (Li et al. (2024)). BRAU-NET++ (Lan et al. (2024)) enhanced feature fusion by element-wise addition between encoder and decoder features, followed by channel and spatial attention. The proposed Skip Connection Channel-Spatial Attention (SCCSA) mechanism improves the model’s ability to highlight important features. However, most existing methods focus on enhancing encoder-side features or unidirectional fusion strategies, while overlooking the importance of explicitly modeling semantic correlations between encoder and decoder features for more effective integration of skip connections.
@@ 2.2 Vision Transformer models
Transformer was initially proposed for natural language processing and opened up new avenues for innovation in computer vision tasks (Dosovitskiy et al. (2020)). The Transformer block allows each element in the input sequence to focus on all other elements by a self-attention mechanism, thus constructing pure Transformer models to effectively adapt to complex image scenes and objects of various sizes over CNNs. For example, the hierarchical Swin Transformer with shifted windows was used as a base block to learn global and distant semantic interactions in SwinUNet (Cao et al. (2022)). Gating mechanisms were added to the axial-attention to learn relative positional coding, to further accurately encode long-range interactions in MedT (Valanarasu et al. (2021)). In MISSFormer (Huang et al. (2021)), efficient self-attention and enhanced mix-FFN were introduced to construct an enhanced Transformer block for aligning features with higher consistency. Inspired by dilated convolutions, a dilated Transformer was proposed to perform global self-attention in a dilated manner in D-Former (Wu et al. (2023)), which expanded the receptive field and reduced computational cost without adding patch blocks. However, compared with CNNs, the pure Transformer model is limited by the learning of local features, which impacts the accurate capture of detailed features, especially in the delicate medical image segmentation task.
@@ 2.3 Hybrid CNN-Transformer models
Figure 1: (a) The architecture of FIF-UNet, (b) Squeeze-and-Excitation Network, (c) UpConv Module. “Conv” represents a standard combination of convolution, batch normalization, and activation.
@@ 3.1 Overall architecture
[…]
In the skip connections, a CSI module is proposed to dynamically recalibrate the feature maps by the designed CIU and SIU, with the objective of obtaining the informative target features. In the decoder network, each decoder stage is constructed based on the CoSE module and UpConv module. The proposed CoSE module aims to enhance the representation of critical features by incorporating the SENet mechanism into CNNs. The UpConv module upsamples the resolution of the CoSE outputs by bilinear interpolation, followed by a convolution layer to refine the up-sampled feature maps, as in Figure 1(c). Instead of predicting the segmentation tasks based only on the last decoder stage, in this work, a MLF module is innovatively proposed to effectively fuse the outputs of the decoder stages to enhance the segmentation details by integrating intra- and inter-class features.
Table 1: Detailed architecture of FIF-UNet. Size: C × H × W(C: channels, H: height, W: width), Res: Residual operation, Bi: Bilinear, CC: Channel Concatenation.
Network Stage Operation Input Size Output Size
Encoder 1 Conv3×3(stride=2) + Conv3×3(stride=1) 3 × 256 × 256 96 × 128 × 128
2 MaxViT Block ×2 96 × 128 × 128 96 × 64 × 64
3 MaxViT Block ×2 96 × 64 × 64 192 × 32 × 32
4 MaxViT Block ×5 192 × 32 × 32 384 × 16 × 16
5 MaxViT Block ×2 384 × 16 × 16 768 × 8 × 8
Skip Connection 2 CSI(CIU + SIU) 96 × 64 × 64 96 × 64 × 64
96 × 64 × 64
3 CSI(CIU + SIU) 192 × 32 × 32 192 × 32 × 32
192 × 32 × 32
4 CSI(CIU + SIU) 384 × 16 × 16 384 × 16 × 16
384 × 16 × 16
Decoder 1 CoSE(Res(Conv3×3(stride=1) + BN + ReLU) + Res(SE)) 96 × 128 × 128 48 × 128 × 128
2 CoSE(Res(Conv3×3(stride=1) + BN + ReLU) + Res(SE)) 96 × 64 × 64 96 × 128 × 128
+ Bi(2×) + Conv3×3(stride=1) + BN + ReLU
3 CoSE(Res(Conv3×3(stride=1) + BN + ReLU) + Res(SE)) 192 × 32 × 32 96 × 64 × 64
+ Bi(2×) + Conv3×3(stride=1) + BN + ReLU
4 CoSE(Res(Conv3×3(stride=1) + BN + ReLU) + Res(SE)) 384 × 16 × 16 192 × 32 × 32
+ Bi(2×) + Conv3×3(stride=1) + BN + ReLU
5 CoSE(Res(Conv3×3(stride=1) + BN + ReLU) + Res(SE)) 768 × 8 × 8 384 × 16 × 16
+ Bi(2×) + Conv3×3(stride=1) + BN + ReLU
Output Module (MLF) Conv1×1(stride=1) + CC 48 × 128 × 128 9 × 256 × 256
+ GroupConv3×3(stride=1) + BN + ReLU 96 × 64 × 64
+ Conv5×5(stride=1) 192 × 32 × 32
384 × 16 × 16
Figure 2: MaxViT architecture. L1, L2, L3, and L4 represent the number of cascaded MaxViT blocks in the four stages, respectively.
@@ 3.2 CSI Module
The CSI module is composed of two sequential sub-units: the Channel Interaction Unit (CIU) and the Spatial Interaction Unit (SIU), as shown in Figure 3. The CSI is applied to the skip connections of the UNet architecture, where it takes as input the encoder features at stage $i$ ($i=2,3,4$) and the decoder features at stage $i+1$.
The purpose is to improve the effectiveness of feature transfer across encoder and decoder levels by enhancing the semantic quality of skip connection features. Specifically:
  • The CIU is to align encoder and decoder features in the channel dimension by learning a set of adaptive channel weights, which is expected to emphasize semantically relevant channels. Since encoder and decoder features are often misaligned due to different levels of abstraction, the channel-wise recalibration also benefits to bridge the semantic gap.
  • The SIU then complements the channel-level refinement by capturing spatial correlations between features, which enables the model to maintain spatial consistency and retain structural details to further support the generation of coherent segmentation outputs.
Finally, the enhanced feature map produced by the CSI module is fed into the decoder stage $i$. It is worth noting that the CSI module can be flexibly integrated into any UNet-like architecture, significantly improving the semantic expressiveness of skip connections and strengthening the decoder’s ability to utilize encoder features effectively.
@@ 3.2.1 CIU
As mentioned before, the inputs of the CIU are the feature maps of the encoder stage $i$ ($X$) and the decoder stage $i+1$ ($Y$). The core idea of the CIU is to adjust the importance of input features ($X$ or $Y$) based on another counterpart feature map ($Y$ or $X$), implemented by parallel paths for the inputs separately (Figure 3(a)).
@@ 3.2.2 SIU
In the right part of the X-shaped path, the sigmoid activation function is performed to project the weight elements to [0,1][0,1][ 0 , 1 ], which is further performed on the SIU inputs to reweight their pixel-wise importance. Finally, the output feature map is formulated by addition operations, as in:
@@
In summary, the sequential combination of CIU and SIU enables the CSI module to jointly optimize semantic relevance and spatial coherence, producing more informative and better-aligned feature representations.